OpenStudy (anonymous):

at what point do the curves r1= <2t-6, 5-t, t^2-4t+5> and r2=< 1-s, 3s-6, -s^2+6s-8> intersect? find the angle of intersection

OpenStudy (anonymous):

if they intersect at some point then the solution to system obtain from corresponding components of curves is the point of intersection . so here are i,j,k components 2t-6=1-s (set i components equal ) 5-t=3s-6 (set j components equal ) solve these equations for s and t and let me know .

OpenStudy (anonymous):

k let see

OpenStudy (anonymous):

for s =-2t+5 s=11/3-t/3

OpenStudy (anonymous):

and t im not sure is t=7/2-s/2 and t=3s+11

OpenStudy (anonymous):

i need their numerical values ok let me show you . rearranging the terms 2t+s=7 ........1 t+3s=11 ....... 2 from equation 1 s=7-2t put in second equation at the place of s t+3(7-2t)=11 t+21-6t=11 -5t=-10 t=2 so s=3 t=2 s=3 is the solution to the system .

OpenStudy (anonymous):

put t=2 and s=3 in any of the give curve equation to get the required point .

OpenStudy (anonymous):

r1 and r2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok r1=<-2,3,1>,r2=<-2,3,1> they look the same r1 and r2

OpenStudy (anonymous):

they should be same it means you find the point of intersection correctly :)

OpenStudy (anonymous):

i find with part 1 of the question?

OpenStudy (anonymous):

i'm finish*

OpenStudy (anonymous):

no now find the angle of intersection .

OpenStudy (anonymous):

i need to use the dot product?

OpenStudy (anonymous):

exactly . first take derivative of r1 and then r2 with respect to t then put t=2 and s=3 to get two vectors . then go for dot product . can you do this now ?

OpenStudy (anonymous):

ok @sami-21

OpenStudy (anonymous):

r'1=<2, -1, 2t-4>=<2,1,0> r'2=< -1, 3, -2s+6>=<-1, 3 0>

OpenStudy (anonymous):

one thing more take the derivative of r2 with respect to s !

OpenStudy (anonymous):

i notice that @sami-21

OpenStudy (anonymous):

<2,1,0>*<-1, 3 0>= 1

OpenStudy (anonymous):

good now take dot product .

OpenStudy (anonymous):

still we find the angle?

OpenStudy (anonymous):

need*

OpenStudy (anonymous):

yes if you have two vector A and B then angle between them is given by divide their dor product by product of the magnitude of each vector and take cos inverse to get angle \[\Large \theta =\cos^{-1} \frac {A.B }{\left| A \right|.\left| B \right| }\]

OpenStudy (anonymous):

50